OpenStudy (anonymous):

Tyrone & elias are painting a house. Together it will take 24 hours to paint. Tyrone alone can paint house in 6 hours less than Elias working alone.How long to the nearest hour would it take each to paint alone

OpenStudy (anonymous):

I set up equation as: (1/x-6)+(1/x)=1/24

OpenStudy (zepp):

Sorry, was away :s And yes that's the right equation to solve the question :)

OpenStudy (anonymous):

Are you not able to solve it? Need help solving it?

OpenStudy (anonymous):

1/(x-6) + 1/x = 1/24 (x + x-6)/(x(x-6)) = 1/24 24(2x - 6) = x(x-6) 48x - 144 = x^2 - 6x x^2 - 54x + 144 = 0

OpenStudy (anonymous):

how do i factor that. i cant find the numbers

OpenStudy (anonymous):

Use quadratic formula. (-b +/- SQRT(b^2 - 4ac))/2a You get: (54 +/- SQRT(2916 - 4*144))/2